# Towards an Accessible, Noninvasive Micronutrient Status Assessment Method: A Comprehensive Review of Existing Techniques

### v1

@@ 3.2.2 Sweat-Based Multiplexed Sensors
Multiplexed analyses combine and analyze data from assays with multiple sensors such as electrocardiogram (ECG), temperature, electrodermal activity (EDA), HRV data and more. Thus, these multiplexed sensors can enable more complex physiologic monitoring and diagnosis. Kim et al. [86] developed an on-body biosensing platform that can collect and analyze cortisol, glucose, and vitamin C in sweat using microfluidics (Fig. 1(c)). The device includes a lateral flow assay for cortisol and fluorometric assays for glucose and vitamin C. Assay results are imaged with a smartphone (with special lenses in the case of fluorometry) and are analyzed to yield quantitative results. There are also electrodes for sweat rate and EDA, with Near-Field Communication (NFC) and Radio Frequency (RF) to power them and communicate results. The focus in the paper is on stress indicators. Field-testing of the device involved subjecting participants to “intensive work periods” (interrupted sleep schedule and caffeine intake) for 7 days then rest (regular schedule) and daily vitamin C intake (for 2 participants) for 14 days [86]. Spikes in vitamin C associated with intake could also be observed. The authors also claim they were able to assess changes in circadian rhythms through cortisol, but no trends in glucose were found.
@@
For a more in-depth review of the nuances of creating wearable multimodal sensors with sweat collection and analysis capabilities, we direct the reader to [187].
Table 1: Assay-Based Methods Using Sweat
@@
Method Form Factor Targets Analytes Evaluation Control Results Notes Source
@@
Fluorometry Wearable patch Chloride, sodium, and zinc Sweat Human subjects Ion chromatography (chloride), AAS (sodium), ICP-MS (zinc) Zinc measurement was the most inaccurate and subject to the most variance Statistical analysis of results is not presented [149]
@@
Colorimetry Wearable patch Vitamin C, calcium, zinc, and iron Sweat Human subjects ICP-MS on diluted sweat samples and supplementation Correlations with ICP-MS of 0.926 for Vit C, 0.743 for calcium, 0.895 for zinc, and 0.963 for iron, time dynamics of measurements after supplementation were in line with those of blood Sensor can also supplement micronutrients transdermally [85]
@@
Fluorometric and lateral flow assays Wearable with smartphone analysis Cortisol, glucose, and vitamin C Sweat Human subjects ELISA for cortisol, controlled stress and diet Cortisol aligned with circadian rhythm changes and had $R^{2}$ of 0.7974 with ELISA, observed spikes in Vit C with intake, no trends in glucose Measurement of sweat rate and EDA, with NFC and RF to power them and communicate results [86]
3.2.3 Smartphone-based Quantitative Assays
@@
Smartphones are increasingly being used as analytical platforms for quantitative assays. Often, the phone is either used to photograph the assay results for quantitative analysis [95, 96, 160, 152, 153, 47, 41, 137], or it communicates with a more specialized and standardized sensing device [103, 97, 176]. Lee et al. [95] demonstrate the use of smartphones to image and quantify vitamin D (calcidiol) levels from an immunoassay. After the sample is deposited on the test and it is incubated for a few hours, the assay is imaged using a custom smartphone accessory. Concentrations can then be quantified from the image by comparing the brightness of the detection region to the reference area. The device was evaluated using three levels of known concentrations that span from deficiency to sufficiency, but not excess (Table 18). Although this range is debated. Results were compared to an ELISA test. Motivated by the asymptomatic nature of Vitamin B12 deficiency, the same group developed a smartphone-based assay method for B12 quantification [96]. The ratio of test to control line intensities extracted from the smartphone image are used as an input to a 4-parameter logistic curve to output an estimated vitamin B12 concentration. 12 human subjects provided whole blood samples from a finger prick, and assay results were compared to an Immulite 2000 immunoassay system. On these samples, the Nutriphone failed to accurately determine B12 levels above 441 pg/mL. Unlike the Immulite, this solution appears to be insensitive to the upper spectrum of B12 insufficiency and sufficiency. In addition, while the authors do not specify the form of vitamin B12 targeted by their device, the reported molecular weight is closest to the non-gold-standard cyanocobalamin. The authors suggest that future work should aim to be more effective at lower limits of detection and better account for interferents in whole blood. Iron, as ferritin, has likewise been targeted by this group [160]. Their assay was evaluated in-lab with known concentrations of ferritin in spiked buffer (n=27) and serum samples (n=12) to optimize performance. Human trials were also conducted and results were compared to the Immulite 2000.
@@
Figure 2: The use of a smartphone-based assay for the quantitative assessment of iron in serum. Used with permission from [152].
@@
Serhan et al. [152] had a similar goal of using a smartphone-based assay to measure total iron in serum (Fig. 2). This paper focuses on total iron instead of ferritin (the clinically-accepted biomarker for imbalance; Table 19) because it is "the most direct metabolite in the [iron] panel" [152]. Total iron can provide valuable insights into iron status, thus it is a worthwhile target even if it is not the gold-standard [118]. This is the first work to consider the risks of excess iron, purposefully designing the assay to be sensitive to both deficiency and excess with a “dynamic range of 50-300 $\mu$g/dL” [152]. 20 finger-prick samples were taken, and assay results were compared to optimized multi-well plate spectrophotometry. Additional analysis found that their approach had a coefficient of variance of 10.5%, compared to 2.2% for the lab tests. One unique form of validation here that did not appear in other work was specificity testing using interferent analytes. The same first author improved their method and developed a new system to measure total iron levels from whole blood, consisting of an iOS smartphone application, a 3D printed sensing chamber, and a vertical flow membrane-based sensor strip [153]. Their new approach filters out the cellular components of whole blood and carries out a colorimetric chelation reaction. A resulting color change is detected by the smartphone device within 5 minutes. The smartphone application’s accuracy and precision were tested against a reference imaging software (ImageJ) for the same colorimetric sensing strip. They compared their iron detection technique to a spectrophotometry-based laboratory test for iron detection on 14 venous blood samples from 9 volunteers and found greater limit of detection (LoD) than the laboratory method (2.2 mcg/dL) The authors discuss future steps of expanding this tool toward measuring total iron binding capacity and saturation levels.
Table 2: Assay-Based Methods Using Smartphones
Method Form Factor Targets Analytes Evaluation Control Results Notes Source
Quantitative immunoassay Smartphone-based Vitamin D (calcidiol) Serum Known sample solutions and human subjects ELISA Errors “at same order” as ELISA Erickson Lab at Cornell; no statistical analysis of results [95]
Quantitative immunoassay Smartphone-based Vitamin B12 (cyanocobalamin) Finger prick blood sample Human Subjects Immulite 2000 Immunoassay Correlation of 0.93 with control,  85% specificity and  60% sensitivity for deficiency detection Erickson Lab at Cornell; used synthetic form of B12; poor accuracy outside of deficiency range [96]
Quantitative immunoassay Smartphone-based or desktop device (TIDBIT) Iron (ferritin) Finger prick blood sample Human Subjects Immulite 2000 Immunoassay Correlation of 0.92 with control, sensitivity of 0.9 for deficiency detection Erickson Lab at Cornell [160]
Quantitative immunoassay Smartphone-based Total iron Finger prick serum blood sample Human Subjects Multi-Well Plate Spectrophotometry $R^{2}$ of 0.98 with control, CoV of 10.5% Conducted specificity testing using interferent analytes; considered toxicity [152]
Quantitative immunoassay Smartphone-based Total iron Finger prick whole blood sample Human Subjects Laboratory developed test: spectrophotometry-based technique Correlation plot with slope of 1.09, R2 of 0.96, and a mean bias of 5.3% Improved on control LoD; manually diluted some samples to represent a low iron concentration [153]
Colorimetric assay Paper-based Total iron Urine Human subjects AAS RSD of 9.5% Urine citric acid was found to interfere with results [47]
Colorimetric assay Paper-based Iron (Fe3+) Serum Spiked, diluted from human subjects Spiked known concentration Error of 3.7% and RSD of 1% LoD of 0.3 mcg/mLl [41]
Colorimetric assay Paper-based Iron (Fe3+) Saliva Spiked pooled commercial saliva Spiked known concentration $R^{2}$ of 0.99 It is debated whether salivary iron reflects circulating status [137]
Opto-electronic immunoassay Card Vitamin A (RBP), Iron (ferritin), and CRP Whole blood Human subjects ELISA CVs of 2.5% for ferritin, 10.8% for RBP, and 3.9% for CRP Cutoff for iron deficiency was insensitive to upper limit; RBP is not the clinically-accepted biomarker for vitamin A status; Device transmits results to a smartphone over NFC [97]
Multiplexed quantitative assay Desktop device (TIDBIT) Vitamin A (RBP), Iron (ferritin), and CRP Serum Human subjects ELISA $R^{2}$ of 0.56 for RBP, 0.92 for ferritin, 0.88 for CRP Erickson Lab at Cornell; ferritin range does not cover anemia; RBP is not the clinically-accepted biomarker for vitamin A status; RMSE for RBP was 21 mcg/mL [103]
Quantitative immunoassay Desktop device (TIDBIT) Vitamin D3 (calcidiol) Serum and whole blood Commercial standards; human subjects Known solutions; LC-MS/MS CoV with standards of 2.63% at 34 ng/mL and 11.2% at 0 ng/mL; $R^{2}$ of 0.91 for serum tests, 0.94 for finger prick tests Erickson lab [176]
@@
Ferreira et al. [47] develop an assay for urine, a biofluid that is less invasive than blood and more representative of status than sweat. Their paper presents a paper-based colorimetric assay for urinary iron and quantification of results from images. The assay itself contains four columns of five sample units each. This design allows for replicate results and outlier exclusion. Blank (water) samples were used to obtain baseline signal intensity for absorbance calculations and calibrate for urine color interference. Calibration curves were determined using iron standards in water and synthetic urine. It was found that the phosphate and citric acid in the synthetic urine significantly interfered with the slope of the calibration curve. The sensor was evaluated using volunteer urine samples (n=26) pre-treated with nitric acid. The results were compared to AAS and found to be similar. Dortez et al. [41] showcase another colorimetric assay for serum iron, quantified by smartphone images. Samples are preprocessed with anti-transferrin to strip iron (Fe3+) ions from the protein transferrin, which are then reduced to Fe2+ before the sample is deposited on the assay. A smartphone and light box were used to capture images of the test area, which are analyzed to yield quantified results. The assay was developed using iron standard solutions and evaluated using diluted serum spiked with Fe3+. It was found that the assay could analyze multiple samples simultaneously, allowing for auto calibration of test samples against a certified reference control. Prakobdi et al. [137] claims to present a low-cost, non-invasive saliva-based screening test for IDA using a nitrocellulose lateral flow system. The method uses a capillary flow-driven microfluidic device and tests iron’s reaction with bathophenanthroline (Bphen) and ferrous (Fe2+) ions to measure iron (Fe3+) levels in spiked saliva. A color change is analyzed, and the results indicate a linear response in the 100-2000 mcg/dL range (falling far above the lower limit for ‘normal’ serum iron results [27]). The study used pooled commercial saliva, and it does not compare measured iron levels with a clinical gold standard for iron status assessment. In addition, we note the general ambiguity of whether saliva levels of iron are an accurate indicator of circulating blood levels (Section 2.2).
@@
The smartphone moves into a supporting role in Lee et al. [97]. The authors implement a paper-based microfluidic immunoassay for vitamin A (as retinol binding protein or RBP), iron (as ferritin), and C-reactive protein (CRP) in a unique, card-like form factor. Most notably, it is able to analyze finger prick whole blood samples with minimal pre-treatment because of built-in plasma separation (similar to [153]). The device includes competitive assays for CRP and RBP and a sandwich assay for ferritin. Sandwich-based assays are more sensitive, while competitive assays are more effective for analytes of larger concentrations. CRP is a useful inclusion since both iron and vitamin A assessment are impacted by inflammation (Table 18). On-device light emitting diodes (LEDs) and photodetectors analyze the assay, sending results and various test metrics to a smartphone app over NFC. The added ability to store this data in a remote server makes the device a powerful tool for population-level screening. The device was evaluated on whole blood samples (n=95), each run 3 times, and compared to ELISA. 84.4% of sample was male, and 6 samples were spiked with ferritin, CRP, and RBP to assess a wider range of concentrations. In assessment, a physiologically-relevant cutoff of 15 mcg/L was set for ferritin deficiency (although this is insensitive to the deficiency upper limit of 30 mcg/L; Table 19). The decision to make this device single-use is perplexing, especially since a reusible version was used for testing. We would also like to note that RBP is not the clinically-accepted biomarker for vitamin A in serum, which should be measured directly instead (Table 18).
@@
In Lu et al. [103], the Erickson lab deviates from smartphone image analysis and moves to a bespoke device. Their new approach achieves simultaneous quantification of vitamin A (as RBP), iron (as ferritin), and CRP on a single test strip using multiple fluorescent markers and immunoassays. The paper proposes a reusable, standalone reader (the TIDBIT). Six blue LEDs with band-pass optical filters (458 nm) are used for fluorescent excitation of the assay tags. Interestingly, quantitative results are only presented to the user if they are within a "physiologically relevant dynamic range" of 2.2-20 $\mu$g/mL for RBP, 12-200 mcg/L for ferritin, and 0.5-10 $\mu$g/mL for CRP [103]. We note that the ferritin range sufficiently covers deficient status, but not anemia (Table 19). 43 human serum samples were purchased from a commercial vendor for testing and compared against ELISA. Evaluation results for RBP were poor, with no explanation of why this may be. This TIDBIT device is applied further in Vemulapati et al. [176], targeting vitamin D (as 25(OH)D3). A “novel elution buffer that separates 25(OH)D3 from its binding protein in situ” enables the assessment of finger prick blood with no pre-treatment [176]. In testing, the test/control ratio was highly correlated to vitamin D concentrations in standard solutions. Commercial serum standards highly correlated with assay results (4-parameter logistic curve), with coefficient of variance (CoV) of 2.63% at 34 ng/mL and 11.2% at 0 ng/mL. Human trials with serum (n=21) and finger prick whole blood (n=6) samples were conducted, and results were compared to LC-MS/MS measurements. The accuracy of deficiency detection for serum but not whole blood was assessed, with an area under the curve (AUC) of 0.836 for deficiency cutoffs of 20 ng/mL and 1 for 12 ng/mL. Only the latter aligns with the general clinical threshold for deficiency (Table 18).
3.2.4 Commercial Products
Commercially-available devices have arose in recent years to provide point of care (POC) testing for some micronutrients [4, 38, 48, 7]. One study has explored the utility of a commercial iCheck FLUORO device to assess vitamin A concentrations in human breast milk (breast milk vitamin A or BMVA) [4]. BMVA is critical since it is the primary source of vitamin A for breastfeeding children. If there is a vitamin A deficiency in breast milk, it is likely to cause developmental issues for a child. The authors collected breast milk samples and socio-demographic and anthropomorphic data from lactating mothers in the Mecha district, Ethiopia (n=104). This region was selected because prior studies applying this device for BMVA assessment recommended further investigation of populations at greater risk of vitamin A deficiency. Concentrations of vitamin A in breast milk were measured by iCheck FLUORO and compared to HPLC. The commercial device was found to overestimate low BMVA concentrations and had a weak overall correlation with HPLC results. Therefore, the paper concluded that studies that assess vitamin A intake among breast-feeding children in developing countries should not assume average BMVA. It is argued that devices like the FLUORO are needed to monitor BMVA status, especially for intervention programs that typically assume average BVMA. Still, they must be “reliable across a range of BMVA concentrations” [4].
Albrecht et al. [7] likewise studied the efficacy of the Quidel Inc Sofia fluorescent immunoassay for serum vitamin D (as calciol). The assay is analyzed by the Sofia Analyzer, a POC device for immunoassay analysis. Because only one sample had a concentration above 100 ng/mL, the authors recommend additional testing for concentrations above 80 ng/mL (well into the range of toxicity; Table 18).
Bloom Diagnostics is a desktop ’lab’ device that analyzes single-use qualitative test strips to quantitatively assess the status of in-vitro (in-body) biomarkers, similar to the Erickson Lab’s TIDBIT [176, 103, 38]. Tests available for Bloom include thyroid-stimulating hormone (TSH), ferritin, CRP, and estimated glomerular filtration rate with cystatin C. While Bloom approaches the goal of accessible nutrition assessment, its assays still require the user to collect a sample themselves. Depending on the test, this could involve a finger prick, coaxing the blood into a collection tube, and depositing it properly onto the assay. VitaScan (from the Erickson lab) is another commercial POC device that tests for iron deficiency, and they validate results against the clinical gold standard for in-body iron measurement [48]. The device is not yet released, but it is planned to assess vitamins B12, D, and A and CRP in the future. The method is still invasive as it utilizes blood obtained from a finger prick and also requires the user to obtain the sample themselves.
Table 3: Commercial Assay-Based Methods

### v2

@@ 3.2.2. Sweat-Based Multiplexed Sensors
Kim et al. (2020) developed an on-body biosensing platform that could collect and analyze cortisol, glucose, and vitamin C in sweat using microfluidics (Fig. 1(iii)). The device included a lateral-flow assay for cortisol and fluorometric assays for glucose and vitamin C. Assay results were imaged with a smartphone (with special lenses in the case of fluorometry) and were analyzed to yield quantitative results. There were also electrodes for sweat rate and EDA. The focus in the paper was on stress indicators. Field-testing of the device for cortisol assessment involved subjecting 2 participants to “intensive work periods” (interrupted sleep schedule and caffeine intake) for 7 days then rest (regular schedule) for 14 days (Kim et al., 2020). Additional tests with 2 different participants for all target biomarkers involved subjecting participants to intensive work followed by a regular schedule and vitamin C supplementation for 14 days. Spikes in vitamin C associated with intake could be observed under these conditions.
@@
For a more in-depth review of the nuances of creating wearable multimodal sensors with sweat collection and analysis capabilities, we direct the reader to (Yokus and Daniele, 2021).
Table 2. Assay-Based Methods Using Sweat
@@
Method Platform Targets Analytes Evaluation Control Results Notes Source
@@
Fluorometry Wearable patch Chloride, sodium, and zinc Sweat Human subjects Ion chromatography (chloride), AAS (sodium), ICP-MS (zinc) Zinc measurement was the most inaccurate and subject to the most variance Statistical analysis of results is not presented (Sekine et al., 2018)
@@
Colorimetry Wearable Vitamin C, calcium, zinc, and iron Sweat Human subjects ICP-MS on diluted sweat samples and supplementation Correlations with ICP-MS of 0.926 for Vit C, 0.743 for calcium, 0.895 for zinc, and 0.963 for iron, time dynamics of measurements after supplementation were in line with those of blood Sensor can also supplement micronutrients transdermally (Kim et al., 2022)
@@
Fluorometric and lateral flow assays Wearable with smartphone analysis Cortisol, glucose, and vitamin C Sweat Human subjects ELISA for cortisol, controlled stress and diet Cortisol aligned with circadian rhythm changes and had $R^{2}$ of 0.7974 with ELISA, observed spikes in Vit C with intake, no trends in glucose Measurement of sweat rate and EDA, with NFC and RF to power them and communicate results (Kim et al., 2020)
3.2.3. Smartphone-based Quantitative Assays
@@
Smartphones are increasingly being used as analytical platforms for quantitative assays. Often, the phone is either used to photograph the assay results for quantitative analysis (Lee et al., 2014, 2016b; Srinivasan et al., 2018; Serhan et al., 2020, 2021; Ferreira et al., 2021; Dortez et al., 2023; Prakobdi et al., 2024), or it communicates with a more specialized and standardized sensing device (Lu et al., 2017; Lee et al., 2016a; Vemulapati et al., 2017). Lee et al. (2014) demonstrated the use of smartphones to image and quantify vitamin D (calcidiol) levels from an immunoassay. After the sample is deposited on the test and incubated for a few hours, the assay was imaged using a custom smartphone accessory. The device was evaluated using three levels of known concentrations that span from deficiency to sufficiency, but not excess, although this range is debated (Table LABEL:tab:_assessment). Results were compared to an ELISA test. The same researchers developed a smartphone-based assay method (dubbed “the Nutriphone”) for B12 quantification (Lee et al., 2016b). Twelve human subjects provided capillary blood samples from a finger prick, and assay results were compared to an Immulite 2000 immunoassay system. On these samples, the Nutriphone failed to accurately determine B12 levels above 441 pg/mL. Unlike the Immulite, this solution appears to be insensitive to the upper spectrum of serum vitamin B12 concentrations (Table LABEL:tab:_assessment). In addition, while the authors do not specify the form of vitamin B12 their device aimed to investigate, the reported molecular weight is closest to the non-optimal reference standard cyanocobalamin. As reported in Table LABEL:tab:_assessment, vitamin B12 has no single optimal reference standard, but it is often confirmed with methylmalonic acid. The authors suggest that future work should aim to be more effective at lower limits of detection and better account for interferents in whole blood. An evolution of the Nutriphone assesses iron, as ferritin (Srinivasan et al., 2018). This assay was evaluated in-lab with known concentrations of ferritin in spiked buffer (n=27) and serum samples (n=12) to optimize performance. Human trials were also conducted (n=20) and results were compared to the Immulite 2000.
@@
Figure 2. The use of a smartphone-based assay for the quantitative assessment of iron in serum. Used with permission from Serhan et al. (2020).
[…]
Top left: a table comparing images of the test area for a blank sample, a serum sample, and iron concentrations ranging from 50 $\mu$g/dL to 1000 $\mu$g/dL. The test area clearly darkens compared to the blank and serum samples as iron concentration increases. Bottom left: a smartphone is shown capturing an image of the test area and displaying the results of the image analysis to the user. Right: the full test setup consists of a smartphone in holder, with the test strip placed into the attachment described in Serhan et al. (2020)
@@
Serhan et al. (2020) had a similar goal of using a smartphone-based assay to measure total iron in serum (Fig. 2). This paper focused on total iron instead of ferritin (the clinically-accepted biomarker for imbalance; Table LABEL:tab:_assessment) because it is ”the most direct metabolite in the [iron] panel” (Serhan et al., 2020). Total iron can provide valuable insights into iron status, thus it is a worthwhile target even if it is not the optimal reference standard (National Institutes of Health, 2023). This is the first work to consider the risks of excess iron, purposefully designing the assay to be sensitive to both deficiency and excess with a “dynamic range of 50-300 $\mu$g/dL” (Serhan et al., 2020). Twenty capillary blood samples were collected via a finger prick, and assay results were compared to optimized multi-well plate spectrophotometry. Additional analysis found that their approach had a coefficient of variance of 10.5%, compared to 2.2% for the lab tests. One unique form of validation here that did not appear in other work was specificity testing using interferent analytes. The preceding method was approved upon with a new system to measure total iron levels from whole blood, consisting of an iOS smartphone application, a 3D printed sensing chamber, and a vertical flow membrane-based sensor strip (Serhan et al., 2021). The smartphone application’s accuracy and precision were tested against a reference imaging software (ImageJ) for the same colorimetric sensing strip. They compared their iron detection technique to a spectrophotometry-based laboratory test for iron detection on 14 venous blood samples from 9 volunteers (7 male) and found greater limit of detection (LoD) than the laboratory method (2.2 $\mu$g/dL) The authors discuss future steps of expanding this tool toward measuring total iron binding capacity and saturation levels.
@@
Ferreira et al. (2021) developed an assay for urine, a biofluid that is less invasive than blood and more representative of status than sweat. Their paper presented a paper-based colorimetric assay for urinary iron and quantification of results from images. The assay itself contained four columns of five sample units each. This design allows for replicate results and outlier exclusion. Blank (water) samples were used to obtain a baseline signal intensity for absorbance calculations and calibrate for urine color interference. Calibration curves were determined using iron standards in water and synthetic urine. It was found that the phosphate and citric acid in the synthetic urine significantly interfered with the slope of the calibration curve. The sensor was evaluated using volunteer urine samples (n=26) pre-treated with nitric acid. The results were compared to AAS and found to be similar. Dortez et al. (2023) showcase another colorimetric assay for serum iron, quantified by smartphone images. The assay was developed using iron standard solutions and evaluated using diluted serum spiked with Fe3+. It was found that the assay could analyze multiple samples simultaneously, allowing for auto calibration of test samples against a certified reference control. Prakobdi et al. (2024) claimed to present a noninvasive saliva-based screening test for IDA using a nitrocellulose lateral flow system to measure iron (Fe3+) levels in spiked saliva. A reaction’s color change was analyzed, and the results indicate a linear response in the 100-2000 $\mu$g/dL range (falling far above the lower limit for ‘normal’ serum iron results (Chad Haldeman-Englert et al., [n. d.])). The study used pooled commercial saliva, and it did not compare measured iron levels with a clinical optimal reference standard for iron status assessment. In addition, we note the general ambiguity of whether saliva levels of iron are an accurate indicator of circulating blood levels (Section 2.2).
@@
The smartphone moved into a supporting role in Lee et al. (2016a). The authors implemented a paper-based microfluidic immunoassay for vitamin A (as retinol binding protein or RBP), iron (as ferritin), and C-reactive protein (CRP) in a unique platform that resembles a flat, portable index-card design. Most notably, it was able to analyze whole blood with minimal pre-treatment because of built-in plasma separation (similar to (Serhan et al., 2021)). The device included competitive assays for CRP and RBP and a sandwich assay for ferritin. Sandwich-based assays are more sensitive, while competitive assays are more effective for analytes of larger concentrations. CRP is a useful inclusion since both iron and vitamin A assessment are impacted by inflammation (Table LABEL:tab:_assessment). On-device light emitting diodes (LEDs) and photodetectors analyzed the assay, sending results and various test metrics to a smartphone app over NFC. The added ability to store this data in a remote server made the device a powerful tool for population-level screening. The device was evaluated on whole blood samples (n=95), each run 3 times, and compared to ELISA. A documented 84.4% of sample was male, and 6 samples were spiked with ferritin, CRP, and RBP to assess a wider range of concentrations. For evaluation, a physiologically-relevant cutoff of 15 $\mu$g/L was set for ferritin deficiency (although this is insensitive to the deficiency upper limit of 30 $\mu$g/L; Table LABEL:tab:_assessment). The decision to make this device single-use is perplexing, especially since a reusable version was used for testing. We would also like to note that RBP is not the clinically-accepted biomarker for vitamin A in serum, which should be measured directly instead (Table LABEL:tab:_assessment).
The approach in Lu et al. (2017) achieved simultaneous quantification of vitamin A (as RBP), iron (as ferritin), and CRP on a single test strip using multiple fluorescent markers and immunoassays. The paper proposed a reusable, standalone reader (the TIDBIT). Interestingly, quantitative results were only presented to the user if they are within a ”physiologically relevant dynamic range” of 2.2-20 $\mu$g/mL for RBP, 12-200 $\mu$g/L for ferritin, and 0.5-10 $\mu$g/mL for CRP (Lu et al., 2017). We note that the ferritin detection range sufficiently covers iron deficient status, but does not extend into the threshold which indicates iron-deficiency anemia (IDA; ¡10 $\mu$g/L; Table LABEL:tab:_assessment). Although a complete blood count (CBC) test that measures hemoglobin is the standard method for diagnosing anemia more generally, an additional ferritin assessment is still useful to determine if the patent suffers from IDA (the most common type of anemia) (National Heart, Lung, and Blood Institute and National Institutes of Health, 2022b). Furthermore, ferritin status can inform the clinician whether iron supplementation is an appropriate treatment for the anemia, as over-supplementation and excess can lead to adverse effects (National Heart, Lung, and Blood Institute and National Institutes of Health, 2022a). To assess the TIDBIT device, forty-three human serum samples were purchased from a commercial vendor for testing and compared against ELISA (Lu et al., 2017). While the R² for the RBP assay was lower than for the other biomarkers, the authors explained that the assay was optimized for high sensitivity and specificity near the diagnostic cutoff for vitamin A deficiency, rather than for precise quantification across the entire physiological range.
@@
This TIDBIT device was applied further in Vemulapati et al. (2017), which examined vitamin D (as 25(OH)D3) via the assessment of capillary blood from a finger prick with no pre-treatment. In testing, the test/control ratio was highly correlated to vitamin D concentrations in standard solutions. Commercial serum standards highly correlated with assay results (4-parameter logistic curve), with coefficient of variance (CoV) of 2.63% at 34 ng/mL and 11.2% at 0 ng/mL. Human trials with serum (n=21) and capillary blood (n=6) samples were conducted, and results were compared to LC-MS/MS measurements. The accuracy of deficiency detection was assessed for serum but not whole blood with an area under the curve (AUC) of 0.836 for deficiency cutoffs of 20 ng/mL and 1 for 12 ng/mL. Only the latter aligns with the general clinical threshold for deficiency (Table LABEL:tab:_assessment).
Table 3. Assay-Based Methods Using Smartphones